The Poseidon Swim Company produces swim trunks. The average selling price for one of their swim trunks is $73.76. The variable cost per unit is $24.80. Poseidon Swim has average fixed costs per year of $8,682. Assume that current level of sales is 304 units. What will be the resulting percentage change in EBIT if they expect units sold to change by -3.8 percent? (You should calculate the degree of operating leverage first).
Answer #1

EBIT when sales is 304 = 304*(73.76 - 24.80) - 8682 = 6201.84

EBIT when sales is 304 *( 1-0.038) = 304*(1-0.038) (73.76 - 24.80) - 8682 = 5636.2541

difference = 565.59

% change = 565.59/6201.84 = 9.12%

DOL = 9.12%/3.8% = 2.40